OpenStudy (anonymous):

Help? I got this wrong.. :( Determine the domain of the function. f(x)=sqrt(1-x)

OpenStudy (anonymous):

\[f(x)=\sqrt{1-x}\]

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

I put x>1..

OpenStudy (anonymous):

Should it be: \[x \le 1\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Dadgummit. -_- Thanks..

OpenStudy (mathmale):

\[f(x)=\sqrt{1-x}\] has restrictions on its domain in that the quantity under the radical sign must always be 0 or greater than 0. So we could write the domain as \[1-x \ge 0\] If you'd add x to both sides, the resulting inequality would be the domain of your function f(x).